HUBBLE SPACE TELESCOPE DAILY REPORT #1003 PERIOD COVERED: 0800 EST Daily Status Report as of 301/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 4110 (The Physics of Massive O-Stars in Different Parent Galaxies) The Faint Object Spectrograph (Blue Side) was used to make observations of giant O class stars in galaxy NGC-346-3 and NGC-346-1. Ultraviolet observations from 1150 A0 to 2300 A0 were made. A couple of recentering events were reported to have occurred during the first observations. Several of the observations were recorded on the science tape recorder. The SHP and UDL were missing for on observation during playback. An HSTAR was written. No other anomalies were reported. 1.2 Completed PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were recorded on science tape recorder #1, and successful played back later. Both observations were reported to be normal exposures. 1.3 Completed FOC/96 5045 (FOC Observations to Monitor Desorption - Cycle 3) The Faint Object Camera was used to make observations of the standard star GSC8842-12 using filter F486N in the F/96 mode to monitor the focus of the HST optics. Two observations were placed on science tape recorder #1, and will be analyzed later. No anomalies were reported. 1.4 Completed FOC/96 5029 (FOC Plate Scale - Cycle 3) The Faint Object camera (f/96) was used to make observations of the object 47 Tuc (aka, NGC-104) to determine the plate scale calibration, and whether it is changing with time. A total of 11 observations were received by the OSS. There was a 23 second recentering event which occurred during on of the observations. No other anomalies were reported. 1.5 Completed FOC/96 4969 (Search for QSOs Suitable for Subsequent Observation of He II) The Faint Object Camera was used to make ultraviolet images of the quasi-stellar object 0215-471 to measure the levels of its ultraviolet flux to see if this object is suitable for subsequent observation with the FOS and GHRS. They are looking for absorption features of Helium II at 304 A0. The observations were executed as scheduled, and no problems were reported. 1.6 Completed HSP/POL 4888 (Remnant Stars In Supernova Remnants - Aug. - Cycle 3 High) The High Speed Photometer was used to make ultraviolet polarimetric observations of super nova remnant SS433 which has emission lines that vary by relativistic amounts in a 162 day cycle. The OSS monitored the observations, and assessed the data quality. No problems were reported. 1.7 Completed WFC 5076 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 2) Routine Earth flat calibration exposures were taken with the Wide Field Camera. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.8 Completed HSP/POL 5039 (GHRS Dark Noise Monitoring - Cycle 3) The High Resolution Spectrograph was used to make dark calibration images. The photoscan mode was used with an integration time of 10 seconds and with seven bins, for a read-out every 70 seconds. There was no carrousel motion. Observations were taken both inside and outside the small science aperture. The observations were executed as scheduled, no anomalies were reported. 1.9 Completed HSP/POL 4889 (Gravitational Lenes Part 1 - Aug. - Cycle 3 High) The High Speed Photometer was used to make observations of quasar QSO2345+00 which is an apparent gravitation lens system. Both photometry and polarmetry observations were made to find its characteristic signature in variability (In the time scale of weeks). The observations were monitored by the OSS, and the data quality was assessed. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Significant HSTARs: None 2.4 Operations Note: The SMGT #14(A) test was initiated Tuesday Oct. 26, and is scheduled to end later today, Thursday, Oct. 28, 1993. All operations are proceeding normally, and no anomalies have been reported. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete PC 4308 (Snapshot Survey for Companions to Planetary Nebula Nuclei - Cycle 3 High ) 3.2 Complete FOS/RD/BL 5051 (FOS Cycle 3 Polarimetric Calibration) 3.3 Complete WFC 4929 (WF/PC Cycle 2 : Improved Dark Calibration) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-